Grade 12th passAlgebra

Find the ratio of
9 m to 27 cm
Find the ratio of
9 m to 27 cm

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

We know that,
1 m = 100 cm
Then,
9 m = 9 × 100 = 900 cm
Now we have to find the ratio,
= 900/27
= 100/3 ... [∵divide both by 9]
So, the required ratio is 100: 3.

Thanks